For each day on the newpaper route you recieve $10.00 plus a fixed amount for each newspaper you deliver. Currently yo are earning $18.00 per day for delivering 80 newspapers. Today you are assigned to deliver 20 additional newspapers per day. What will be your new daily earnings?
A.) 10.00
B.)18.00
C.) 19.00
D.)20.00
E.) 22.50

Found 2 solutions by rfer, jim_thompson5910:
Answer by rfer(16322)

8/80=0.10
100*0.10+10=$20

Answer by jim_thompson5910(35256)

"For each day on the newpaper route you recieve $10.00 plus a fixed amount for each newspaper you deliver."

So you earn 10 + k*x dollars where k is the fixed number you earn per paper and x is the number of papers.

-------------------------------------------------------

"Currently you are earning $18.00 per day for delivering 80 newspapers. "

which means 10+k*x = 10+k*80 = 10+80k is the amount of money made. That amount of money made is $18, so

10+80k = 18

80k = 18-10

80k = 8

k = 8/80

k = 1/10

k = 0.10

So you earn $0.10 or 10 cents per paper.

-------------------------------------------------------

Now you are delivering 20 additional papers. So the total delivered is 80+20 = 100 papers

10+k*x = 10+0.10*100 = 10+10 = 20

The final answer is $20.00